Phys 1408 cal based

Phys 1408 cal based
media%2F9c0%2F9c08cc78-b513-4ce6-a6a5-12
Answer #1

converting time into appropriate units.

2.0 years = 2.0*365*24*60 = 1051200.0

So the number of time her heart beats is 69.0 * 1051200.0 = 72532800.0= 7.25 × 101

2.00 years = 2.00*365*24*60 = 1051200.00

So the number of times her heart beats is 69.0 * 1051200.00 = 72532800.0= 7.25 × 101

2.000 years = 2.000*365*24*60 = 1051200.000

so 69.0*1051200.000 = 72532800.0= 7.25 × 101

You can notice that all of the answer have only one zero. Because while doing the calculations the answer should have least no of significant numbers.